OpenStudy (anonymous):

does someone know how to find (√(3)+i)^(1/3)?

OpenStudy (luigi0210):

\[\sqrt{3+i}^(\frac{ 1 }{ 3 })\] Is this what you mean?

OpenStudy (anonymous):

\[\sqrt{3}+i ^{1/3}\]